# Data-driven Radiative Magnetohydrodynamics Simulations with the MURaM Code: the Emergence of Active Region 11158 and the X2.2 Flare

### v1

@@ 3.4.2 Evolution of the flare ribbons and Coronal Magnetic Reconnection
Figure 6: Comparison of the flare ribbons in the observation and simulation. The top row displays the evolution of flare ribbons captured by the AIA 1600 channel. The middle row shows the flare ribbon in the simulation, which is manifested by the emission measure in the lowermost 200 grid points at all temperatures. The bottom row illustrates magnetic field lines that are closely related to flare ribbons. The grayscale images are $B_{z}$ in the photosphere of the simulation, overlaid by a horizontal slice at 2.5 Mm height showing plasma temperatures, as indicated by the color table. Only temperatures higher than 10 MK are displayed, whereas lower values are transparent. Two groups of magnetic field lines are displayed: one group is calculated from seed points that are randomly distributed in regions with bright ribbons with a bias to high density and the other group has a bias to high temperature. The color of the fieldlines represents the plasma temperature, as indicated by the same color table of the slice.
@@
Finally, it is intriguing to examine the extent to which the flare ribbons in the model may resemble the observation of this flare and what is the magnetic configuration that directly causes the evolution of the ribbon. The AIA 1600Å channel images covering the change in the flare ribbons from the impulsive to the gradual phases are presented in the top row of Figure 6, which is compared with the total emission measure at all temperatures in the lower atmosphere, as shown in Figure 6 (e) – (h). The bottom row presents a top-down view of the magnetic field in a similar field of view.
@@
In the impulsive phase shown in the first column in Figure 6, both the observation and simulation show bright ribbons along the central PIL and a brightening on the right side of the N2 spot. During the peak, while the brightest emission remains in the active region core, the ribbons extend to the center of the P1 spot, with bright patches above strong flux concentrations. The gradual phase in the third column demonstrates that the ribbons in both the observation and model start to fade and exhibit separating motion in the N–S direction. Finally, as time progresses, the two main ribbons become clearly more separated with significantly reduced brightness.
@@
The latter two snapshots presented in Panels (k) and (l) illustrate, in closer view than those presented in Figure 2, the formation of a corridor of post-flare loops. As the large flux rope rises and tiggers reconnection between the large scale envelope field, as shown in Figure 2 (g), longer and higher loops with apparently weak shear form, whose footpoints map the expanded flare ribbon seen in the emission features.
4 Summary and Discussion
@@
The complete three-stage model demonstrates the emergence of the magnetic field from the photosphere to the corona and a major solar eruption that releases $10^{32}$ erg magnetic energy and yields a soft X-ray flux on a similar order of magnitude as the real event. In the large-scale atmosphere, the eruption drives bulk mass ejections and a dome-like shock front; meanwhile, the coronal eruption deliver a sufficiently strong impact to the outer rim of the active region that penetrates untill the chromosphere and creates a fast-propagating Moreton wave. The released magnetic energy is deposited in the chromosphere by strong downward conduction fluxes that are greater than $10^{11}$ erg/cm2/s. Local pressure peaks drive chromospheric condensation and evaporation. The temporal evolution of the bright flare ribbons in the simulation, which maps the development of reconnections in the coronal magnetic field, is consistent with the observation of this flare.
@@
Observations have suggested that the seminal structure of this eruption was a flux rope along the highly sheared PIL between the N2 and P1 spots. Despite of the different magnetic initial conditions and boundary driving methods, data-driven simulations of AR 11158 have consistently reproduced such a magnetic structure (Inoue et al., 2014; Hayashi et al., 2018; Afanasyev et al., 2023; Fan et al., 2024). In our model, we use a simple approach of Cheung & DeRosa (2012). Although the outer part of the active region exhibits some artificial twist at the large scale due to the constant $\Omega$ and horizontal size boundary, the choice of $\Omega=1.0\times 10^{-1}$ s-1 is able to build an appropriate flux rope that creates an eruption of a similar strength as the real flare.
@@
The plasma dynamics in flare loops have been thoroughly investigated in classical hydrodynamics calculations (Fisher et al., 1985). Idealized 3D MHD models suggest that the location of the flare ribbons reflect the cross sections of magnetic separatrixes in the lower atmosphere (e.g., Janvier et al., 2014; Savcheva et al., 2015). The ingredients of these models must be combined in 3D radiative MHD simulations to yield flare ribbons that can be compared with observations, as performed by Cheung et al. (2019) in an MURaM simulation inspired by observations.
@@
Furthermore, reproducing the observed ribbons in this particular flare is more challenging. Inoue et al. (2014) used magnetic properties to project the location of flare ribbons in the model, which appeared to be similar to the observed ribbons. The models presented by Afanasyev et al. (2023) and Fan et al. (2024) stand for the state-of-art of data-driven models of this active region thus far in the literature. The former illustrated the location of flare ribbons by a temperature proxy instead of a direction observable, because of some missing ingredients in the coronal energy balance. The latter improved the energy equation; however, unfortunately, the bottom boundary has a fixed chromospheric temperature and density, which limited investigations of flare ribbons and the chromospheric impact of the coronal shock. Therefore, to the best of our knowledge, the work presented in this paper may be the first model that self-consistently describes the evolution of flare ribbons, which resembles a real flare in a one-to-one fashion.
@@
In addition to studying this particular active region, which has been performed rather thoroughly, the simulations presented in this paper could also be useful for investigating, in general, the gradual evolution of a flux rope before the eruption and the trigger of the onset. The data-driven MURaM code can be of cause applied in general to other observed active regions, as long as a time series of the vertical magnetic field is available. It is also interesting to test in the further how our simulations perform when they are driven by the electric field used in other models Kazachenko et al. (2014); Fisher et al. (2020) or a more sophisticated characteristic boundary developed by Tarr et al. (2024).
@@
F.C. is supported by National Science Foundation of China No. 12422308 and No. 12373054, and by the National Key R&D Program of China under grant 2021YFA1600504. This work benefits from discussions during the ISSI workgroup “Data-driven 3D Modeling of Evolving and Eruptive Solar Active Region Coronae”. The visualization shown in Figures 1, 2, and 6 are created by VAPOR (Li et al., 2019). {contribution}F.C. conducted the numerical simulations, performed the analysis and visualization, and wrote the texts.

### v2

@@ III.4.2 Evolution of the flare ribbons and Coronal Magnetic Reconnection
Figure 6: Comparison of the flare ribbons in the observation and simulation. The top row displays the evolution of flare ribbons captured by the AIA 1600 channel. The middle row shows the flare ribbon in the simulation, which is manifested by the emission measure in the lowermost 200 grid points at all temperatures. The bottom row presents magnetic field lines that are closely related to flare ribbons. The grayscale images are $B_{z}$ in the photosphere of the simulation, overlaid by a horizontal slice at 2.5 Mm height showing plasma temperatures, as indicated by the color table. Only temperatures higher than 10 MK are displayed, whereas lower values are transparent. Two groups of magnetic field lines are displayed: one group is calculated from seed points that are randomly distributed in regions with bright ribbons with a bias to high density and the other group has a bias to high temperature. The color of the fieldlines represents the plasma temperature, as indicated by the same color table of the slice.
@@
Finally, it is intriguing to examine the extent to which the flare ribbons in the model may resemble the observation of this flare and the magnetic configuration that directly causes the evolution of the ribbon. The AIA 1600Å channel images covering the change in the flare ribbons from the impulsive to the gradual phases are presented in the top row of Figure 6. We note that the structures along the y-axis in Panels (a) and (b) are caused by saturation. The observation is compared with the total emission measure at all temperatures in the lower atmosphere, as shown in Figure 6 (e)–(h). The bottom row presents a top-down view of the magnetic field in a similar field of view.
@@
In the impulsive phase shown in the first column in Figure 6, both the observation and simulation show bright ribbons along the central PIL and a brightening on the right side of the N2 spot. During the peak, while the brightest emission remains in the active region core, the ribbons extend to the center of the P1 spot, with bright patches above strong flux concentrations. The gradual phase in the third column demonstrates that the ribbons in both the observation and model start to fade and exhibit separating motion in the N–S direction. Finally, as time progresses, the two main ribbons clearly become more separated with significantly reduced brightness.
@@
The latter two snapshots presented in Panels (k) and (l) illustrate, in closer view than those presented in Figure 2, the formation of a corridor of post-flare loops. As the large flux rope rises and triggers reconnection between the large scale envelope field, as shown in Figure 2 (g), longer and higher loops with apparently weak shear form, whose footpoints map the expanded flare ribbon seen in the emission features.
IV Summary and Discussion
@@
The complete three-stage model demonstrates the emergence of the magnetic field from the photosphere to the corona and a major solar eruption that releases $10^{32}$ erg magnetic energy and yields a soft X-ray flux on a similar order of magnitude as the real event. In the large-scale atmosphere, the eruption drives bulk mass ejections and a dome-like shock front; meanwhile, the coronal eruption delivers a sufficiently strong impact on the outer rim of the active region, which penetrates to the chromosphere and creates a fast-propagating Moreton wave. The released magnetic energy is deposited in the chromosphere by strong downward conduction fluxes that are greater than $10^{11}$ erg/cm2/s. Local pressure peaks drive chromospheric condensation and evaporation. The temporal evolution of the bright flare ribbons in the simulation, which maps the development of reconnections in the coronal magnetic field, is consistent with the observation of this flare.
@@
Observations have suggested that the seminal structure of this eruption was a flux rope along the highly sheared PIL between the N2 and P1 spots. Despite the different magnetic initial conditions and boundary driving methods, data-driven simulations of AR 11158 have consistently reproduced such a magnetic structure (S. Inoue et al., 2014; K. Hayashi et al., 2018; A. N. Afanasyev et al., 2023; Y. Fan et al., 2024). In our model, we use a simple approach of M. C. M. Cheung & M. L. DeRosa (2012). Although the outer part of the active region exhibits some artificial twist at the large scale due to the constant $\Omega$ and horizontal size boundary, the choice of $\Omega=1.0\times 10^{-1}$ s-1 is able to build an appropriate flux rope that creates an eruption of a similar strength as the real flare.
@@
The plasma dynamics in flare loops have been thoroughly investigated via classical hydrodynamics calculations (G. H. Fisher et al., 1985). Idealized 3D MHD models suggest that the locations of flare ribbons reflect the cross sections of magnetic separatrices in the lower atmosphere (e.g., M. Janvier et al., 2014; A. Savcheva et al., 2015). The ingredients of these models must be combined in 3D radiative MHD simulations to yield flare ribbons that can be compared with observations, as performed by M. C. M. Cheung et al. (2019) in an MURaM simulation inspired by observations.
@@
Furthermore, reproducing the observed ribbons in this particular flare is more challenging. S. Inoue et al. (2014) used magnetic properties to project the location of flare ribbons in the model, which appeared to be similar to the observed ribbons. The models presented by A. N. Afanasyev et al. (2023) and Y. Fan et al. (2024) stand for the state of the art of data-driven models of this active region thus far in the literature. The former employed a simplified plasma energy equation but was able to illustrate the location of flare ribbons via a temperature proxy. This proxy is different from observed chromospheric flare ribbons because, in that model, the lower boundary actually represents the coronal base, and temperature evolution is affected by a boundary of fixed values. The latter improved the energy equation and reproduced coronal EUV observations; however, unfortunately, the bottom boundary has a fixed chromospheric temperature and density, which still limits investigations of flare ribbons and the chromospheric impact of the coronal shock. In comparison, our model has a stratified solar atmosphere starting from the photosphere, which is consistent with the layer of the observed magnetic field, and the chromospheric density and temperature dynamically evolve in response to the energy injection from the corona. Therefore, to the best of our knowledge, the work presented in this paper may be the first model that self-consistently describes the evolution of flare ribbons, which resembles a real flare in a one-to-one fashion.
@@
In addition to studying this particular active region, which has been performed rather thoroughly, the simulations presented in this paper could also be useful for investigating, in general, the gradual evolution of a flux rope before eruption and the trigger of onset. The data-driven MURaM code can be of cause applied in general to other observed active regions, as long as a time series of the vertical magnetic field is available. It is also interesting to test in the further how our simulations perform when they are driven by the electric field used in other models M. D. Kazachenko et al. (2014); G. H. Fisher et al. (2020) or a more sophisticated characteristic boundary developed by L. A. Tarr et al. (2024).
@@
F.C. is supported by National Science Foundation of China No. 12422308 and No. 12373054, and by the National Key R&D Program of China under grant 2021YFA1600504. This work benefits from discussions during the ISSI workgroup “Data-driven 3D Modeling of Evolving and Eruptive Solar Active Region Coronae”. The visualizations shown in Figures 1, 2, and 6 are created by VAPOR (S. Li et al., 2019).
{contribution}
F.C. conducted the numerical simulations, performed the analysis and visualization, and wrote the texts.
Appendix A Control experiments on the temporal and spatial scaling factors in zero-$\beta$ models
Figure 7: Comparison of magnetic energy in zero-$\beta$ runs with different speed-up and volume scaling factors. The upper panel presents the evolution of the total ($E_{\rm tot}$, solid lines) and free ($E_{\rm free}$, dashed lines) magnetic energy over time. The lower panel shows the energy change rate, i.e., the time derivative of the magnetic energy between February 12 and February 14, following the line colors noted in the upper panel. Four cases are compared. The time and energy of each case are multiplied by its corresponding speed-up factor $f_{\rm sp}$ and the volume scaling factor $f_{\rm V}$, respectively. The vertical dotted line indicates the time of the real X2.2 flare.
Figure 8: Comparison of the magnetic energy in the zero-$\beta$ and the MHD evo. runs shown in Figure 1. The time and energy of the zero-$\beta$ run are multiplied by $f_{\rm sp}$ and $f_{\rm V}$, respectively.
As described in Section II, the evolution of the observed magnetic field that drives the zero-$\beta$ model is accelerated by a factor $f_{\rm sp}$, and the domain of the zero-$\beta$ model represents a smaller region than the actual active region. The robustness of the results is substantiated by further control experiments.
[…]
We compare the total and free magnetic energy of these runs in Figure 7. The speed and volume factors are listed in the legend for each case. All the cases show a quantitatively consistent trend following the emergence of the active region over an equivalent time of 4.5 solar days. A factor that causes the differences between the curves of difference cases is some weak activities that occur before February 15. However, the slopes of the curves, which describe the increasing rate of the magnetic energy, are consistent across all the runs. This is demonstrated in the lower panel of Figure 7, where the time derivatives of the curves in the upper panel are shown following the same line styles.
The consistency of the results from runs with and without a scaling in time and space can also be examined by comparing the zero-$\beta$ run and the MHD evo. run that are already shown in Figure 1. In Figure 8, we compare the magnetic energy of the zero-$\beta$ run, which is scaled by the corresponding factors $f_{\rm sp}=12$ and $f_{\rm V}=1.875^{3}$, and that of the MHD evo. run, which has the same domain size as the actual active region and is driven by the magnetic field at the original cadence.
The small difference in the total magnetic energy curves is caused by that the potential part of the initial condition of the MHD evo. run is directly calculated from the observed magnetic field (see Section II), which gives the MHD evo. run a slightly higher starting point. In comparison, the difference between the $E_{\rm free}$ curves is approximately 1% or smaller.
The curves of magnetic energy in both run exhibit a consistent trend for more than 2 hours until $t\approx-3000$ s. The gentle decreasing trend in the zero-$\beta$ run curves (dotted lines) indicates that the eruption is about to occur, whereas the eruption in the MHD evo. run some minutes later ($t\approx 200$ s).